OpenStudy (anonymous):

How do I extract common factors from this equation? 6(2x+5)^2(3x-4)-6(2x+5)^3/(3x-4)3

OpenStudy (anonymous):

My coursework shows this... 6(2x+5)^2(3x-4-2x-5)/(3x-4)^3 I just don't understand how they get it??

OpenStudy (ash2326):

We have \[6(2x+5)^2(3x-4)-\frac{6(2x+5)^3}{(3x-4)^3}\] We can notice the common term (2x+5) in both the terms \[6\underline{(2x+5)^2}(3x-4)-\frac{6\underline {(2x+5)^3}}{(3x-4)^3}\] We can take (2x+5)^2 common from both the terms so we get \[(2x+5)^2(6(3x-4)-\frac{6(2x+5)}{(3x-4)^3})\] We can take 6 also common from both the terms, we get \[6(2x+5)^2((3x-4)-\frac{(2x+5)}{(3x-4)^3})\]

OpenStudy (anonymous):

Good explanation. Thank you.